# Unified Linear Parametric Map Modeling and Perception-aware Trajectory Planning for Mobile Robotics

### v1

@@ Abstract
Autonomous navigation in mobile robots, reliant on perception and planning, faces major hurdles in large-scale, complex environments. These include heavy computational burdens for mapping, sensor occlusion failures for UAVs, and traversal challenges on irregular terrain for UGVs, all compounded by a lack of perception-aware strategies. To address these challenges, we introduce Random Mapping and Random Projection (RMRP). This method constructs a lightweight linear parametric map by first mapping data to a high-dimensional space, followed by a sparse random projection for dimensionality reduction. Our novel Residual Energy Preservation Theorem provides theoretical guarantees for this process, ensuring critical geometric properties are preserved. Based on this map, we propose the RPATR (Robust Perception-Aware Trajectory Planner) framework. For UAVs, our method unifies grid and Euclidean Signed Distance Field (ESDF) maps. The front-end uses an analytical occupancy gradient to refine initial paths for safety and smoothness, while the back-end uses a closed-form ESDF for trajectory optimization. Leveraging the trained RMRP model’s generalization, the planner predicts unobserved areas for proactive navigation. For UGVs, the model characterizes terrain and provides closed-form gradients, enabling online planning to circumvent large holes. Validated in diverse scenarios, our framework demonstrates superior mapping performance in time, memory, and accuracy, and enables computationally efficient, safe navigation for high-speed UAVs and UGVs. The code will be released to foster community collaboration.
@@ I Introduction
2) This theorem provides end-to-end guarantees for applying SRP to classification and regression tasks, preserving model stability and generalization. It yields a precise dimension-selection formula, $k=\Theta(p\log(p/\delta)/\varepsilon^{2})$, and reduces computational cost from $O(dk)$ to $O(s\,k)$.
@@ IV-B Feasibility and Existence Analysis
Claim: For any target vector T and any arbitrarily small positive real number $\epsilon>0$, there exists a natural number P, such that when the random mapping dimension M satisfies $M>P$, at least one linear model parameter $\beta$ exists that satisfies the condition $||\beta^{T}S-T||<\epsilon$.
@@ V Random Projection
    Sparse Random Projection (SRP). In order to reduce computational complexity without sacrificing essential information, the high-dimensional vector $\mathbf{x}$ is mapped to a lower-dimensional space via SRP. Let $\mathbf{R}\in\mathbb{R}^{k\times d}$ be a $\{0,\pm 1\}$-valued random matrix constructed according to Achlioptas’ database-friendly scheme, with density parameter $1-s$. Such a construction ensures that the cost of computing $\mathbf{R}\,\mathbf{x}$ scales in proportion to the number of nonzero entries of $\mathbf{x}$, yielding a time complexity of $O\bigl{(}\mathrm{nnz}(\mathbf{x})\bigr{)}$ for matrix–vector multiplication in high-dimensional sparse settings [57, 58]. The projected vector is then given by
@@ V-A Projection Decomposition of the Residual Vector
$\mathbf{e}_{\mathrm{proj}}$ $:=\;\underbrace{\mathbf{R}\mathbf{x}}_{\text{projected sample}}-\;\underbrace{P_{\mathbf{R}\mathcal{S}}\,\mathbf{R}\mathbf{x}}_{\text{projection onto }\mathbf{R}\mathcal{S}}$ (7)
@@
  • Geometric Interpretation: The vector $\mathbf{e}_{\mathrm{proj}}$ is exactly the orthogonal projection of $\mathbf{R}\,\mathbf{x}_{\perp}$ onto the orthogonal complement of $\mathbf{R}\mathcal{S}$ in $\mathbb{R}^{k}$. Its dimension is $k-p$, analogous to the classical least‐squares residual lying in $\mathbb{R}^{n-p}$ [62].
@@ V-B Sin $\Theta$ Angle Leakage Bound
[…]
@@ V-C Sparse Hanson–Wright Concentration Inequality
Beyond the subspace rotation error quantified by (9), one must ensure that the $\ell_{2}$-norm $\bigl{\|}\mathbf{R}\,\mathbf{x}_{\perp}\bigr{\|}$ remains tightly concentrated around $\|\mathbf{x}_{\perp}\|$ in a single SRP instance. To this end, we employ a sparse Hanson–Wright inequality, which generalizes the classical Hanson–Wright bound to the setting of $\{0,\pm 1\}$-valued matrices with density $1-s$ [65, 66]. In particular, for any fixed vector $\mathbf{x}_{\perp}$, if $\mathbf{R}$ is drawn according to the sparse random model of density $1-s$, then
@@
The proof relies on sub-Gaussian tail bounds for “striped” matrix structures and accounts for the sparsity parameter $1-s$, yielding an error bound of the same order as in the dense JL setting [65].
@@ V-D Main Residual Energy Theorem
[…]
@@
satisfies $\Pr[\mathcal{E}_{\mathrm{tot}}]\geq 1-\delta$. The reason one can require each sub-event to have failure probability at most $\delta/3$ is that, in what follows, one chooses the projection dimension $k$ sufficiently large so that each of the three conditions—$\varepsilon$-subspace embedding, Davis–Kahan angle bound, and sparse Hanson–Wright concentration—fails with probability at most $\delta/3$. Hence the overall failure probability does not exceed $\delta$.
@@
[…]
@@
The Residual Energy Preservation Theorem above combines the $\varepsilon$-subspace embedding property with classical linear-model error bounds to fill a gap in random mapping theory: it gives a quantitative analysis of how the energy of the component orthogonal to $\mathcal{S}$ is preserved after projection. Specifically, when the projection dimension satisfies (5), inequalities (12)–(13) hold with $C_{2}=1+\sqrt{2}\,(1+\varepsilon)\approx 3.9$ (for $\varepsilon\leq 0.3$). This result provides clear guidance for choosing dimensionality‐reduction parameters: one can ensure that the projected residual remains within an $O(\varepsilon)$ factor of the original residual energy, and one can reduce the failure probability $\delta$ by increasing $k$ appropriately. In large‐scale sparse feature settings, SRP not only reduces the matrix–vector multiplication cost from $O(d\,k)$ to $O(s\,k)$, but also significantly accelerates computation while preserving approximation accuracy, thereby guaranteeing model stability and generalization after projection.
@@ VI-C Trajectory Optimization for UAVs
$+\lambda_{d}\sum_{\{x,y,z\}}\left\{\sum_{i=p_{b}-1}^{N-p_{b}}\mathbf{F}(v_{\max}^{2},\tilde{q}_{i,\mu}^{2})+\sum_{i=p_{b}-2}^{N-p_{b}}\mathbf{F}(a_{\max}^{2},\tilde{q}_{i,\mu}^{2})\right\}$
@@
$\left.F_{c}(\mathbf{D(\mathbf{Q}_{i})})=\left\{\begin{matrix}(\mathbf{D(\mathbf{Q}_{i})}-S_{f})^{2}&\mathbf{D(\mathbf{Q}_{i})}\leq S_{f}\\ 0&\mathbf{D(\mathbf{Q}_{i})}>S_{f}\end{matrix}\right.\right.$ (21)
@@
Input: RMRP model $(\eta,W,b,g)$; start/goal $x_{s}$, $x_{g}$; weights $\lambda_{g}$, $\lambda_{d}$; iterations $N$; step size $\alpha$
@@ 3 $P^{*}\leftarrow P_{0}$;
5for $k\leftarrow 1$ to $N$ do
6       $\nabla J\leftarrow$ Grad$(P^{*},P_{0},\text{RMRP},\lambda_{g},\lambda_{d})$;
@@ VII-A Analytical Derivation of the Terrain Elevation Gradient
$=\sum_{i=1}^{D}\beta_{i}\nabla_{x}\left[\sin\left(W_{i}^{T}x+b_{i}\right)\right]$ (26)
@@ VII-B Terrain Penalty Term for Trajectory Optimization
$=\sum_{i=p_{b}}^{N-p_{b}}\left\|W\left(\beta\odot\cos\left(W^{T}Q_{i}+b\right)\right)\right\|^{2}$ (33)
@@ VII-C Gradient of the Terrain Penalty Term
$=W\left(\beta\odot\left[-\sin\left(\mathbf{y}_{i}\right)\right]\right)W^{T}$ (41)
@@
$=2W\left(\beta\odot\left[-\sin\left(\mathbf{y}_{i}\right)\right]\right)W^{T}\mathbf{g}\left(Q_{i}\right)$ (43)
@@
×$\quad\times W{\begin{pmatrix}\beta_{1}\cos(y_{i1})&0&\cdots&0\\ 0&\beta_{2}\cos(y_{i2})&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\beta_{D}\cos(y_{iD})\end{pmatrix}}$
@@ IX-A Derivation of the Residual Energy Preservation Theorem
[…]

### v2

@@ Abstract
Autonomous navigation in mobile robots, reliant on perception and planning, faces major hurdles in large-scale, complex environments. These include heavy computational burdens for mapping, sensor occlusion failures for UAVs, and traversal challenges on irregular terrain for UGVs, all compounded by a lack of perception-aware strategies. To address these challenges, we introduce Random Mapping and Random Projection (RMRP). This method constructs a lightweight linear parametric map by first mapping data to a high-dimensional space, followed by a sparse random projection for dimensionality reduction. Our novel Residual Energy Preservation Theorem provides theoretical guarantees for this process, ensuring critical geometric properties are preserved. Based on this map, we propose the RPATR (Robust Perception-Aware Trajectory Planner) framework. For UAVs, our method unifies grid and Euclidean Signed Distance Field (ESDF) maps. The front-end uses an analytical occupancy gradient to refine initial paths for safety and smoothness, while the back-end uses a closed-form ESDF for trajectory optimization. Leveraging the trained RMRP model’s generalization, the planner predicts unobserved areas for proactive navigation. For UGVs, the model characterizes terrain and provides closed-form gradients, enabling online planning to circumvent large holes. Validated in diverse scenarios, our framework demonstrates superior mapping performance in time, memory, and accuracy, and enables computationally efficient, safe navigation for high-speed UAVs and UGVs.
@@ I Introduction
2) This theorem provides end-to-end guarantees for applying SRP to classification and regression tasks, preserving model stability and generalization. It yields a precise dimension-selection formula, $k=\Theta(p\log(p/\delta)/\varepsilon^{2})$, and reduces computational cost from $O(dk)$ to O(sk)O(s\,k).
@@ IV-B Feasibility and Existence Analysis
Claim: For any target vector T and any arbitrarily small positive real number $\epsilon>0$, there exists a natural number P, such that when the random mapping dimension M satisfies $M>P$, at least one linear model parameter $\beta$ exists that satisfies the condition $\beta^{T}S-T||<\epsilon$.
@@ V Random Projection
    Sparse Random Projection (SRP). In order to reduce computational complexity without sacrificing essential information, the high-dimensional vector $\mathbf{x}$ is mapped to a lower-dimensional space via SRP. Let $\mathbf{R}\in\mathbb{R}^{k\times d}$ be a $\{0,\pm 1\}$-valued random matrix constructed according to Achlioptas’ database-friendly scheme, with density parameter 1s1-s. Such a construction ensures that the cost of computing $\mathbf{R}\,\mathbf{x}$ scales in proportion to the number of nonzero entries of $\mathbf{x}$, yielding a time complexity of $O\bigl{(}\mathrm{nnz}(\mathbf{x})\bigr{)}$ for matrix–vector multiplication in high-dimensional sparse settings [57, 58]. The projected vector is then given by
@@ V-A Projection Decomposition of the Residual Vector
$\mathbf{e}_{\mathrm{proj}}$ $=\;\underbrace{\mathbf{R}\mathbf{x}}_{\text{projected sample}}-\;\underbrace{P_{\mathbf{R}\mathcal{S}}\,\mathbf{R}\mathbf{x}}_{\text{projection onto }\mathbf{R}\mathcal{S}}$ (7)
@@
  • Geometric Interpretation: The vector $\mathbf{e}_{\mathrm{proj}}$ is exactly the orthogonal projection of $\mathbf{R}\,\mathbf{x}_{\perp}$ onto the orthogonal complement of $\mathbf{R}\mathcal{S}$ in $\mathbb{R}^{k}$. Its dimension is kpk-p, analogous to the classical least‐squares residual lying in $\mathbb{R}^{n-p}$ [62].
@@ V-B Sin $\Theta$ Angle Leakage Bound
$\leq\sqrt{2}\,(1+\varepsilon)\,\varepsilon\,\|\mathbf{x}_{\perp}\|\quad\text{(by $\varepsilon$-embedding)}$
@@ V-C Sparse Hanson–Wright Concentration Inequality
Beyond the subspace rotation error quantified by (9), one must ensure that the $\ell_{2}$-norm $\bigl{\|}\mathbf{R}\,\mathbf{x}_{\perp}\bigr{\|}$ remains tightly concentrated around $\|\mathbf{x}_{\perp}\|$ in a single SRP instance. To this end, we employ a sparse Hanson–Wright inequality, which generalizes the classical Hanson–Wright bound to the setting of $\{0,\pm 1\}$-valued matrices with density 1s1-s [65, 66]. In particular, for any fixed vector $\mathbf{x}_{\perp}$, if $\mathbf{R}$ is drawn according to the sparse random model of density 1s1-s, then
@@
The proof relies on sub-Gaussian tail bounds for “striped” matrix structures and accounts for the sparsity parameter 1s1-s, yielding an error bound of the same order as in the dense JL setting [65].
@@ V-D Main Residual Energy Theorem
$\Pr\bigl{[}\mathcal{E}_{\mathrm{len}}\bigr{]}\;,\;\Pr\bigl{[}\mathcal{E}_{\mathrm{ang}}\bigr{]}\;,\;\Pr\bigl{[}\mathcal{E}_{\mathrm{hw}}\bigr{]}\;\geq\;1-\frac{\delta}{3}.$ (11)
@@
[…]
@@
$\bigl{|}\|\mathbf{e}_{\mathrm{proj}}\|^{2}-\|\mathbf{x}_{\perp}\|^{2}\bigr{|}\;\leq\;2\,C_{2}\,\varepsilon\,\|\mathbf{x}_{\perp}\|^{2}.$ (13)
@@
The Residual Energy Preservation Theorem above combines the $\varepsilon$-subspace embedding property with classical linear-model error bounds to fill a gap in random mapping theory: it gives a quantitative analysis of how the energy of the component orthogonal to $\mathcal{S}$ is preserved after projection. Specifically, when the projection dimension satisfies (5), inequalities (12)–(13) hold with $C_{2}=1+\sqrt{2}\,(1+\varepsilon)\approx 3.9$ (for $\varepsilon\leq 0.3$). This result provides clear guidance for choosing dimensionality‐reduction parameters: one can ensure that the projected residual remains within an $O(\varepsilon)$ factor of the original residual energy, and one can reduce the failure probability $\delta$ by increasing $k$ appropriately. In large‐scale sparse feature settings, SRP not only reduces the matrix–vector multiplication cost from O(dk)O(d\,k) to O(sk)O(s\,k), but also significantly accelerates computation while preserving approximation accuracy, thereby guaranteeing model stability and generalization after projection.
@@ VI-C Trajectory Optimization for UAVs
$\hskip-10.00002pt+\lambda_{d}\sum_{\{x,y,z\}}\left\{\sum_{i=p_{b}-1}^{N-p_{b}}\mathbf{F}(v_{\max}^{2},\tilde{q}_{i,\mu}^{2})+\sum_{i=p_{b}-2}^{N-p_{b}}\mathbf{F}(a_{\max}^{2},\tilde{q}_{i,\mu}^{2})\right\}$
@@
[…]
@@
Input: RMRP model $(\eta,W,b,g)$; start/goal $x_{s}$, $x_{g}$; weights $\lambda_{g}$, $\lambda_{d}$; iterations NN; step size $\alpha$
@@ 3 $P^{*}\leftarrow P_{0}$;
5for $k\leftarrow 1$ to NN do
[…]
@@ VII-A Analytical Derivation of the Terrain Elevation Gradient
$\sum_{i=1}^{D}\beta_{i}\nabla_{x}\left[\sin\left(W_{i}^{T}x+b_{i}\right)\right]$ (26)
@@ VII-B Terrain Penalty Term for Trajectory Optimization
$\sum_{i=p_{b}}^{N-p_{b}}\left\|W\left(\beta\odot\cos\left(W^{T}Q_{i}+b\right)\right)\right\|^{2}$ (33)
@@ VII-C Gradient of the Terrain Penalty Term
$W\left(\beta\odot\left[-\sin\left(\mathbf{y}_{i}\right)\right]\right)W^{T}$ (41)
@@
$2W\left(\beta\odot\left[-\sin\left(\mathbf{y}_{i}\right)\right]\right)W^{T}\mathbf{g}\left(Q_{i}\right)$ (43)
@@
$\quad\times W{\begin{pmatrix}\beta_{1}\cos(y_{i1})&0&\cdots&0\\ 0&\beta_{2}\cos(y_{i2})&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\beta_{D}\cos(y_{iD})\end{pmatrix}}$
@@ IX-A Derivation of the Residual Energy Preservation Theorem
$\bigl{|}\|\mathbf{e}_{\mathrm{proj}}\|^{2}-\|\mathbf{x}_{\perp}\|^{2}\bigr{|}\;\leq\;2\,C_{2}\,\varepsilon\,\|\mathbf{x}_{\perp}\|^{2}.$ (52)